# Electric dipole polarizability constraints on neutron skin and symmetry energy

### v1

@@ 3 Introduction
As detailed below, all relevant theoretical models predict a strong correlation of $L$ with two experimentally accessible quantities, viz. the neutron skin thickness and the dipole polarizability. The connection is illustrated in Fig. 2. The density distributions of neutrons $\rho_{n}(r)$ and protons $\rho_{p}(r)$ in the ground state of can be determined from the condition of minimum energy. They approximately have the shape of Fermi distributions as illustrated in Fig. 2 for a nucleus with $N\gg Z$. The mean square radius of neutrons, $R_{n}=\left<r^{2}\right>^{1/2}_{n}$ is slightly larger than that of protons $R_{p}=\left<r^{2}\right>^{1/2}_{p}$. The difference between the two, $r_{\rm skin}=R_{n}-R_{p}$ is defined as the neutron skin thickness.
@@
The dipole polarizability (DP) of nuclei can be obtained from measurements of the photoabsorption cross sections. A connection between DP, neutron skin thickness and parameters of the symmetry energy can only be made through models. Such calculations are presently based either on density functional theory (DFT) [12] or ab initio coupled cluster calculations [39] using interactions derived from chiral effective field theory ($\chi$EFT) [28]. Both types of models predict a strong correlation between the magnitudes of dipole polarizability, $r_{\rm skin}$, and $L$.
While several experimental techniques to measure the DP are discussed, the present review mainly focuses on recent progress using relativistic Coulomb excitation in forward-angle proton scattering at energies of several hundred MeV [129]. One advantage of this method is consistent results across the neutron separation energy, while most of the other experimental techniques are limited to either the energy region below or above.
@@ 4 Relation between dipole polarizability, neutron skin thickness and symmetry energy
The dipole polarizability $\alpha_{\rm D}$ is related to the reduced $B(E1)$ transition strengths and the photoabsorption cross sections $\sigma_{\rm abs}$ by
$\alpha_{\mathrm{D}}=\frac{\hbar c}{2\pi^{2}}\int_{0}^{\infty}\frac{\sigma_{\mathrm{abs}}}{E_{\rm x}^{2}}{\rm d}E_{\rm x}=\frac{8\pi}{9}\int_{0}^{\infty}\frac{\rm B(E1)}{E_{\rm x}}{\rm d}E_{\rm x}.$ (4)
@@ 4.2 Connections in ab initio models
Predictions of $\alpha_{\rm D}$ and correlations with proton and neutron radii based on $\chi$EFT interactions have been obtained from calculations based on a coupled-cluster expansion of the wave functions [39] combined with the Lorentz-integral-transform approach to extract the $E1$ strength [8]. An example of such calculations for 48Ca [38] is presented in Fig. 5(B), where the correlation of $r_{\rm p}$ with $r_{\rm n}$, $r_{\rm skin}$ and $\alpha_{\rm D}$ is displayed together with representative examples of DFT predictions. While the DFT results predict neutron skin values ranging from 0.16 fm to 0.22 fm, the ab initio results based on a set of interactions from Refs. [44, 27] consistently predict rather small values varying from 0.12 to 0.15 fm.
A major difference between the two theoretical approaches lies in the predicted relation between proton and neutron radius. The DFT predictions of $r_{\rm p}$ are approximately constant, most likely because the charge radius of 48Ca is in all cases part of the data set used to fix the model parameters. The ab initio calculations, on the other hand, predict a linear correlation, leading to the approximate constancy of the neutron skin. The absolute value of $\alpha_{\rm D}$ in the ab initio models shows a larger variation compared to the DFT calculations, but can be well described by a linear correlation similar to $r_{\rm n}$. As discussed in the following, these correlations permit for an extraction of constraints on the range of symmetry energy parameters based on the successful description of experimentally measured polarizabilities and charge radii. This type of calculation has been limited so far to closed-(sub)shell nuclei. For recent attempts of an extension to open-shell nuclei see Refs. [17, 19].
@@ 5.1 Experimental Methods
In this section, we discuss the experimental methods to extract the $B(E1)$ distribution in nuclei and the DP. It is technically difficult to directly measure the DP of nuclei as the response to a static electric field although there exist exceptional cases of very light nuclei; see e.g. the works studying the deviation of elastic scattering cross section from Rutherford scattering [102, 69]. Instead, $B(E1)$ or $\sigma_{abs}$ are measured and integrated to determine the DP by Eq. (4).
Some of the experimental methods discussed below are restricted in the accessible excitation energy range, i.e., the techniques are applicable below or above neutron emission threshold ($S_{n}$) only. Thus, the role of contributions to the DP below $S_{n}$, from the IVGDR and from the energy region above the IVGDR are discussed in more detail.
@@ 5.1.1 Photoneutron measurement
Neutron emission is the dominant decay process after photoexcitation for a nucleus as heavy as ${}^{208}{\rm Pb}$ since charged particle decays are strongly suppressed by the Coulomb barrier and the $\gamma$ decay branch is as low as 1-2% [11]. Thus, the $(\gamma,xn)$ cross sections in heavy nuclei can be compared with total photoabsorption cross sections. The ${}^{208}{\rm Pb}(\gamma,xn)$ cross sections are plotted in Fig. 6. The data were taken at LLNL (light blue open circles [42] and blue solid circles [14]) and at Saclay (red half-filled circles [124]). The results from the two laboratories show clear discrepancies, which is also true for some other nuclei. Kawano et al. [51] reported that ”in general, the Saclay $(\gamma,1n)$ cross sections are larger than the Livermore data, whereas the Saclay $(\gamma,2n)$ cross sections are smaller than the corresponding Livermore data”.
@@ 5.1.2 Total photoabsorption
Total photon absorption, $(\gamma,{\rm abs})$, was studied by applying transmission measurements. In this method, the attenuation of photons in a thick target was measured as a function of the photon energy for extraction of the photoabsorption cross sections. At the Mainz electron accelerator, a narrow photon beam was produced by Bremsstrahlung of an electron beam. The average photon flux was $10^{9}$ photons/MeV at 20 MeV. Two identical Compton spectrometers monitored the photon flux before and after a natural abundance target with a thickness of 40-200 cm [5]. The dominant atomic photoabsorption cross sections needed to be subtracted. A high-resolution transmission measurement at AIST was reported for ${}^{28}{\rm Si}$ using a HPGe detector [41]. Recently an experimental setup for photon transmission measurement is under operation at the photon tagger NEPTUN [106] at the S-DALINAC accelerator in Darmstadt.
@@ 5.2 Decomposition of $E1$ and $M1$ contributions
Figure 7: (A) Top: Double differential cross sections of the 120Sn($\vec{p},\vec{p}^{\prime}$) reaction (black squares) and decomposition into non-spinflip (red diamonds) and spinflip (blue circles) parts. The green solid line shows the cross sections due to excitation of the ISGQR. Bottom: Total spin transfer as defined in Ref. [111]. Figure taken from Ref. [43]. (B) Top: Spectra of the 40Ca($p,p^{\prime}$) reaction at $E_{\rm p}=295$ MeV and different scattering angles. Bottom: Example of the decomposition of the spectrum at $\Theta_{\rm lab}=0.4^{\circ}$ into contributions of $L>1$ multipoles (orange), continuum background (green) and $E1$ (red). Figure taken from Ref. [31].
@@
An example of the MDA analysis is presented in Fig. 7(B) for 40Ca [31]. Spectra at different scattering angles are displayed in the upper panel demonstrating strongly forward-peaked cross sections in the energy region of the IVGDR expected for Coulomb excitation. The lower panel shows the partial contributions to the cross sectios at the most forward angle measured resulting from the MDA: $E1$ (red), multipoles $L>1$ (orange) and nuclear background from quasifree scattering (green). Note that $M1$ strength was neglected for 40Ca because it is concentrated in a single state [37]. A comparison of the two independent methods was made in studies of 96Mo [76] and 208Pb [113] and good correspondence of the resulting $E1$ and $M1$ cross sections was found. The results shown in Fig. 7(A) have also been confirmed in a MDA analysis [9]. Since the polarization transfer measurements require secondary scattering, statistics are limited. Thus, the $E1/M1$ decomposition was restricted to MDA in most $(p,p^{\prime})$ experiments.
@@ 5.4 Contributions from the PDR
All particle-emission coincidence experiments accessing the $E1$ strength are limited to the excitation region above the lowest particle separation threshold. Experimental evidence has accumulated that in nuclei with neutron excess significant $E1$ strength – often concentrated in a resonance-like structure commonly termed PDR – can be found below [105, 18]. Its contribution to the DP can be significant because of the inverse energy weighting, cf. Eq. (4). As examples, they amount to about 10% in 58Ni [19] and 8-13 % in the stable Sn isotopes [10].
@@
The origin of the low-energy $E1$ strength in nuclei with neutron excess is a topic of current debate. It has been suggested to arise from an oscillation of the excess neutrons forming a skin against the (approximately) isospin-saturated core [87, 60]. If true, its strength should be related to the neutron skin thickness and in turn to the parameters of the symmetry energy [52, 21, 15]. However, a recent study of the Sn isotope chain for mass numbers 111-124 casts doubts on such a picture. The correlation between neutron excess and neutron skin thickness in Sn isotopes has been experimentally demonstrated with different methods [57], but based on combined data from Oslo [74, 72, 71] and ($p,p^{\prime}$) experiments [9] only a minor fraction of the photoabsorption cross section (expressed as fraction of the TRK sum rule) can be related to the PDR [73]. A decomposition into the tail of the IVGDR and two resonance-like structures is shown in Fig. 9(B) [75]. The contribution interpreted as PDR is much smaller than those of the IVGDR and the prominent structure around 8 MeV. These findings rather point to an interpretation of the PDR as low-energy part of a toroidal $E1$ mode [98, 127]. At present, understanding the nature of the PDR remains an open problem. It is clear, however, that DFT predictions restricted to 1p-1h excitations cannot be used to estimate the low-energy $E1$ strength for cases where data are unavailable [75].
@@ 5.5 Contributions from high excitation energies
At excitation energies beyond the giant resonance region photonuclear cross sections typically contribute a few percent only to the DP. However, for precision results they need to be considered. Data up the pion threshold have been measured for a few cases, viz. natCa [5], natSn [65] and 208Pb [124, 107]. They show approximately constant cross sections as function of excitation energy and were considered for the extraction of the DP from (${}^{\prime})$ experiments [113, 43] neglecting an isotopic dependence. The dominant excitation mechanism in this energy regime is the quasideuteron effect [66]. It has been pointed out by Roca-Maza et al. [101] that these contributions are not included in model calculations based on DFT and should thus be removed in comparison to theoretical predictions. For heavy nuclei they can be estimated using Ref. [23], while in light nuclei they are negligible in the energy range covered by the models [16, 31, 19].
@@ 6 Extracting neutron skin thickness and symmetry energy properties from dipole polarizability data
In this section we discuss constraints derived from the comparison between model predictions and experimental studies of the DP.
@@ 6.1 Constraints based on density functional theory
The predictions are displayed as filled ellipses which represent the $1\sigma$ error as defined in Ref. [96]. The DD functional performs rather well. The other models tend to slightly overestimate the experimental mean values of both 40Ca and 48Ca, but their $1\sigma$ error ellipses do overlap with the experimental bands, except for PC. In all cases, the $\alpha_{\rm D}$ values for both nuclei are highly correlated. The dashed ellipses show the effect of additionally including the experimental $\alpha_{\rm D}$ value of 208Pb [113] in the fit yielding functionals denoted ”-alpha”. This improves the agreement with experiment and shrinks the error ellipsoids. The models incorporate a span of symmetry energy parameters $J=30-35$ $(30-32)$ MeV and $L=32-82$ $(35-52)$ MeV for the calculations excluding (including) the 208Pb data point.
@@
Roca-Maza et al. [101] combined the experimental DP data for 68Ni [103], 120Sn [43] and 208Pb [113] to test a large variety of density functionals. Since the DFT calculatons do not include contributions from the quasi-deuteron process dominating the photoabsorption cross sections above the energy region of the IVGDR, these had to be removed for a comparison [101]. Figure 10(D) presents correlation plots between the experimental results and theoretical predictions from a wide range of DFT interactions. Only a handful (marked in red) is capable of simultaneously describing all three data points. Based on this reduced set, systematic predictions of $\alpha_{\rm D}$ for other masses, $r_{\rm skin}$ and the symmetry energy parameters could be derived. The experimental results for 40,48Ca discussed above are fairly well described by these predictions.

### v3

@@ 3 Introduction
As detailed below, all relevant theoretical models predict a strong correlation of $L$ with two experimentally accessible quantities, viz. the neutron skin thickness and the dipole polarizability. The connection is illustrated in Fig. 2. The density distributions of neutrons $\rho_{n}(r)$ and protons $\rho_{p}(r)$ in the ground state of can be determined from the condition of minimum energy. They approximately have the shape of Fermi distributions as illustrated in Fig. 2 for a nucleus with $N\gg Z$. The mean square radius of neutrons, $R_{n}=\left<r^{2}\right>^{1/2}_{n}$ is slightly larger than that of protons $R_{p}=\left<r^{2}\right>^{1/2}_{p}$. The difference between the two, $r_{skin}=R_{n}-R_{p}$ is defined as the neutron skin thickness.
@@
The dipole polarizability (DP) of nuclei can be obtained from measurements of the photoabsorption cross sections. A connection between DP, neutron skin thickness and parameters of the symmetry energy can only be made through models. Such calculations are presently based either on density functional theory (DFT) [13] or ab initio coupled cluster calculations [40] using interactions derived from chiral effective field theory ($\chi$EFT) [29]. Both types of models predict a strong correlation between the magnitudes of dipole polarizability, $r_{\rm skin}$, and $L$. The considerable experimental challenges of direct measurements of the neutron skin and the model dependencies of methods extracting the neutron skin from the difference of mass and charge radius [117] call for an alternative experimental observable. Since properties of the symmetry energy cannot be extracted directly from experiments but require theory input, measurement of the dipole polarizability provides independent constraints.
While several experimental techniques to measure the DP are discussed, the present review mainly focuses on recent progress using relativistic Coulomb excitation in forward-angle proton scattering at energies of several hundred MeV [131]. One advantage of this method is consistent results across the neutron separation energy, while many of the other experimental techniques are limited to either the energy region below or above. Even more important, measurements of the $E1$ strength with relativistic Coulomb excitation can be extended to exotic nuclei at rare isotope beam facilities like RIKEN, FRIB and GSI/FAIR. Such experiments are performed in inverse kinematics, where the virtual photon flux can be boosted by using a high-$Z$ target and efficient setups with almost $4\pi$ solid angle coverage for detection of neutron emission above [3] and $\gamma$ emission below neutron threshold [134, 104]. In combination with the large cross sections this will permit access to nuclei with extremely large neutron excess much closer to the properties of neutron-rich matter relevant to the physics of neutron stars. Besides the PUMA project [7] aiming at the neutron skin thickness in unstable nuclei using antiproton annihilation, dipole polarizability measurements with relativistic Coulomb exitation are probably the only experimental probe promising insight into properties of the symmetry energy over a wide range of neutron-to-proton ratios.
The paper is organized as follows. Section 4 discusses how information on the neutron skin thickness and symmetry energy can be inferred from model calculations based on DFT (Sect. 4.1) and ab initio methods (4.2). Section 5 is devoted to experimental issues. A short discussion of the available techniques in Sec. 5.1 is followed by a description of methods to disentangle electric and magnetic contributions to the DP in Sec. 5.2. The relevance of experimental information in the energy region of the IVGDR as well as below the neutron threshold and above the IVGDR is compared in Secs. 5.3-5.5. The comparison of experimental and theoretical results (Sec. 6) for a range of nuclei from 40Ca to 208Pb and constraints on neutron skin thickness and paramaters of the symmetry energy extracted thereof are presented in Sec. 6.1 for DFT and Sec. 6.2 for ab initio approaches. Section 6.3 focuses on the difficulties to simultaneously describe results of parity-violating elastic electron scattering and DP experiments with present-day models. Finally, Sec. 6.4 discusses systematics of the DP and the role of volume and surface contributions to the symmetry energy. A summary and an outlook are given in Sec. 7.
@@ 4 Relation between dipole polarizability, neutron skin thickness and symmetry energy
In this section we discuss how information on the neutron skin thickness and parameters of the symmetry energy can be inferred from the comparison of the experimental dipole polarizability to theoretical predictions. At the moment, there are two classes of models either based on DFT or an ab initio coupled cluster approach. Since isovector observables are not well constrained in DFT, quantitative predictions of the DP can vary considerably. However, one can establish a robust correlation between the parameters $J$ and $L$ of the symmetry energy through $\alpha_{D}$. With ab initio based models one aims at an absolute prediction of $\alpha_{D}$ and the underlying symmetry energy parameters of the interaction can be used to calculate the EOS.
The dipole polarizability $\alpha_{D}$ is related to the reduced $B(E1)$ transition strengths and the photoabsorption cross sections $\sigma_{\rm abs}$ by
$\alpha_{D}=\frac{\hbar c}{2\pi^{2}}\int_{0}^{\infty}\frac{\sigma_{abs}}{E_{x}^{2}}{\rm d}E_{x}=\frac{8\pi}{9}\int_{0}^{\infty}\frac{B(E1)}{E_{x}}{\rm d}E_{x}.$ (4)
@@ 4.2 Connections in ab initio models
Predictions of $\alpha_{D}$ and correlations with proton and neutron radii based on $\chi$EFT interactions have been obtained from calculations based on a coupled-cluster expansion of the wave functions [40] combined with the Lorentz-integral-transform approach to extract the $E1$ strength [9]. An example of such calculations for 48Ca [39] is presented in Fig. 5(B), where the correlation of $r_{p}$ with $r_{n}$, $r_{kin}$ and $\alpha_{D}$ is displayed together with representative examples of DFT predictions. While the DFT results predict neutron skin values ranging from 0.16 fm to 0.22 fm, the ab initio results based on a set of interactions from Refs. [45, 28] consistently predict rather small values varying from 0.12 to 0.15 fm.
A major difference between the two theoretical approaches lies in the predicted relation between proton and neutron radius. The DFT predictions of $r_{p}$ are approximately constant, most likely because the charge radius of 48Ca is in all cases part of the data set used to fix the model parameters. The ab initio calculations, on the other hand, predict a linear correlation, leading to the approximate constancy of the neutron skin. The absolute value of $\alpha_{D}$ in the ab initio models shows a larger variation compared to the DFT calculations, but can be well described by a linear correlation similar to $r_{n}$. As discussed in the following, these correlations permit for an extraction of constraints on the range of symmetry energy parameters based on the successful description of experimentally measured polarizabilities and charge radii. This type of calculation has been limited so far to closed-(sub)shell nuclei. For recent attempts of an extension to open-shell nuclei see Refs. [18, 20].
@@ 5 Dipole polarizability from experiment
In this section, we discuss the experimental methods to extract the $B(E1)$ distribution in nuclei and the DP. It is technically difficult to directly measure the DP of nuclei as the response to a static electric field although there exist exceptional cases of very light nuclei; see e.g. the works studying the deviation of elastic scattering cross section from Rutherford scattering [103, 70]. Instead, $B(E1)$ or $\sigma_{abs}$ distributions are measured and integrated to determine the DP by Eq. (4). Some of the experimental methods discussed below are restricted in the accessible excitation energy range, i.e., the techniques are applicable below or above neutron emission threshold ($S_{n}$) only. Thus, the role of contributions to the DP below $S_{n}$, from the IVGDR and from the energy region above the IVGDR are discussed in more detail. Both $E1$ and $M1$ transitions are excited and possible ways of their distinction are briefly presented.
@@ 5.1.1 Photoneutron measurement
Neutron emission is the dominant decay process after photoexcitation for a nucleus as heavy as ${}^{208}{\rm Pb}$ since charged particle decays are strongly suppressed by the Coulomb barrier and the $\gamma$ decay branch is as low as 1-2% [12]. Thus, the $(\gamma,xn)$ cross sections in heavy nuclei can be compared with total photoabsorption cross sections. The ${}^{208}{\rm Pb}(\gamma,xn)$ cross sections are plotted in Fig. 6. The data were taken at LLNL (light blue open circles [43] and blue solid circles [15]) and at Saclay (red half-filled circles [126]). The results from the two laboratories show clear discrepancies, which is also true for some other nuclei. Kawano et al. [52] reported that "in general, the Saclay $(\gamma,n)$ cross sections are larger than the Livermore data, whereas the Saclay $(\gamma,2n)$ cross sections are smaller than the corresponding Livermore data".
@@ 5.1.2 Total photoabsorption
Total photon absorption was studied by applying transmission measurements. In this method, the attenuation of photons in a thick target was measured as a function of the photon energy for extraction of the photoabsorption cross sections. At the Mainz electron accelerator, a narrow photon beam was produced by Bremsstrahlung of an electron beam. The average photon flux was $10^{9}$ photons/MeV at 20 MeV. Two identical Compton spectrometers monitored the photon flux before and after a natural abundance target with a thickness of 40-200 cm [5]. The dominant atomic photoabsorption cross sections needed to be subtracted. A high-resolution transmission measurement at AIST was reported for ${}^{28}{\rm Si}$ using a HPGe detector [42]. Recently an experimental setup for photon transmission measurement is under operation at the photon tagger NEPTUN [107] at the S-DALINAC accelerator in Darmstadt.
@@ 5.2 Decomposition of $E1$ and $M1$ contributions
Figure 7: (A) Top: Double differential cross sections of the 120Sn($\vec{p}$,$\vec{p}^{\prime}$) reaction (black squares) and decomposition into non-spinflip (red diamonds) and spinflip (blue circles) parts. The green solid line shows the cross sections due to excitation of the ISGQR. Bottom: Total spin transfer as defined in Ref. [112]. Figure taken from Ref. [44]. (B) Top: Spectra of the 40Ca$(p,p^{\prime})$ reaction at $E_{p}=295$ MeV and different scattering angles. Bottom: Example of the decomposition of the spectrum at $\Theta_{lab}=0.4^{\circ}$ into contributions of $L>1$ multipoles (orange), continuum background (green) and $E1$ (red). Figure taken from Ref. [32].
@@
An example of the MDA analysis is presented in Fig. 7(B) for 40Ca [32]. Spectra at different scattering angles are displayed in the upper panel demonstrating strongly forward-peaked cross sections in the energy region of the IVGDR expected for Coulomb excitation. The lower panel shows the partial contributions to the cross sectios at the most forward angle measured resulting from the MDA: $E1$ (red), multipoles $L>1$ (orange) and nuclear background from quasifree scattering (green). Note that $M1$ strength was neglected for 40Ca because it is concentrated in a single state [38]. A comparison of the two independent methods was made in studies of 96Mo [77], 120Sn [44] and 208Pb [114] and good correspondence of the resulting $E1$ and $M1$ cross sections was found. The results shown in Fig. 7(A) have also been confirmed in a MDA analysis [10]. Since the polarization transfer measurements require secondary scattering, statistics are limited. Thus, in most $(p,p^{\prime})$ experiments the $E1/M1$ decomposition was restricted to MDA.
@@ 5.4 Contributions from the PDR
All particle-emission coincidence experiments accessing the $E1$ strength are limited to the excitation region above the lowest particle separation threshold. Experimental evidence has accumulated that in nuclei with neutron excess significant $E1$ strength – often concentrated in a resonance-like structure commonly termed PDR – can be found below [106, 19]. Low-energy $E1$ strength is also found in lighter nuclei with $N\approx Z$. Its contribution to the DP can be significant because of the inverse energy weighting, cf. Eq. (4). As examples, they amount to about 10% in 58Ni [20] and 8-13 % in the stable Sn isotopes [11].
@@
The origin of the low-energy $E1$ strength in nuclei with neutron excess is a topic of current debate. It has been suggested to arise from an oscillation of the excess neutrons forming a skin against the (approximately) isospin-saturated core [88, 61]. If true, its strength should be related to the neutron skin thickness and in turn to the parameters of the symmetry energy [53, 22, 16]. However, a recent study of the Sn isotope chain for mass numbers 111-124 casts doubts on such a picture. The correlation between neutron excess and neutron skin thickness in Sn isotopes has been experimentally demonstrated with different methods [58], but based on combined data from Oslo [75, 73, 72] and ($p,p^{\prime}$) experiments [10] only a minor fraction of the photoabsorption cross section (expressed as fraction of the TRK sum rule) can be related to the PDR [74]. A decomposition into the tail of the IVGDR and two resonance-like structures is shown in Fig. 9(B) [76]. The contribution interpreted as PDR is much smaller than those of the IVGDR and the prominent structure around 8 MeV. These findings rather point to an interpretation of the PDR as low-energy part of a toroidal $E1$ mode [99, 129]. At present, understanding the nature of the PDR remains an open problem. It is clear, however, that DFT predictions restricted to 1p-1h excitations cannot reliably estimate the low-energy $E1$ strength distribution for cases where data are unavailable [76].
@@ 5.5 Contributions from high excitation energies
At excitation energies beyond the giant resonance region photonuclear cross sections typically contribute a few percent only to the DP. However, for precision results they need to be considered. Data up the pion threshold have been measured for a few cases, viz. natCa [5], natSn [66] and 208Pb [126, 108]. They show approximately constant cross sections as function of excitation energy and were considered for the extraction of the DP from $(p,p^{\prime})$ experiments [114, 44] neglecting an isotopic dependence. The dominant excitation mechanism in this energy regime is the quasideuteron effect [67]. It has been pointed out by Roca-Maza et al. [102] that these contributions are not included in model calculations based on DFT and should thus be removed in comparison to theoretical predictions. For heavy nuclei they can be estimated using Ref. [24], while in light nuclei they are negligible in the energy range covered by the models [17, 32, 20].
@@ 6 Extracting neutron skin thickness and symmetry energy properties from dipole polarizability data
In this section we discuss constraints on the neutron skin thickness and symmetry energy properties derived from the comparison between model predictions and experimental studies of the DP. These refer to specific nuclei like 40Ca, 48Ca and 208Pb but also systematic isotopic trends or a global mass dependence. The difficulties of presently available models to simultaneously account for measured polarizabilities and asymmetries in parity-violating elastic electron scattering are illuminated.
@@ 6.1 Constraints based on density functional theory
The predictions are displayed as filled ellipses which represent the $1\sigma$ error as defined in Ref. [97]. The DD functional performs rather well. The other models tend to slightly overestimate the experimental mean values of both 40Ca and 48Ca, but their $1\sigma$ error ellipses do overlap with the experimental bands, except for PC. In all cases, the $\alpha_{D}$ values for both nuclei are highly correlated. The dashed ellipses show the effect of additionally including the experimental $\alpha_{D}$ value of 208Pb [114] in the fit yielding functionals denoted "-alpha". This improves the agreement with experiment and shrinks the error ellipsoids. The models incorporate a span of symmetry energy parameters $J=30-35$ $(30-32)$ MeV and $L=32-82$ $(35-52)$ MeV for the calculations excluding (including) the 208Pb data point.
@@
Roca-Maza et al. [102] combined the experimental DP data for 68Ni [104], 120Sn [44] and 208Pb [114] to test a large variety of density functionals. Since the DFT calculatons do not include contributions from the quasi-deuteron process dominating the photoabsorption cross sections above the energy region of the IVGDR, these had to be removed for a comparison [102]. Figure 10(D) presents correlation plots between the experimental results and theoretical predictions from a wide range of DFT interactions. Only a handful (marked in red) is capable of simultaneously describing all three data points. Based on this reduced set, systematic predictions of $\alpha_{D}$ for other masses, $r_{skin}$ and the symmetry energy parameters could be derived. The experimental results for 40,48Ca discussed above are fairly well described by these predictions.